# Geometry-Aware Edge Pooling for Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) have shown significant success for graph-based tasks. Motivated by the prevalence of large datasets in real-world applications, pooling layers are crucial components of GNNs. By reducing the size of input graphs, pooling enables faster training and potentially better generalisation. However, existing pooling operations often optimise for the learning task at the expense of fundamental graph structures and interpretability. This leads to unreliable performance across varying dataset types, downstream tasks and pooling ratios. Addressing these concerns, we propose novel graph pooling layers for structure-aware pooling via edge collapses. Our methods leverage diffusion geometry and iteratively reduce a graph’s size while preserving both its metric structure and structural diversity. We guide pooling using magnitude, an isometry-invariant diversity measure, which permits us to control the fidelity of the pooling process. Further, we use the spread of a metric space as a faster and more stable alternative ensuring computational efficiency. Empirical results demonstrate that our methods (i) achieve superior performance compared to alternative pooling layers across a range of diverse graph classification tasks, (ii) preserve key spectral properties of the input graphs, and (iii) retain high accuracy across varying pooling ratios.
@@ 1 Introduction
Graph pooling layers are important components of successful GNN architectures. They are implemented alongside convolutional layers to reduce the size of graph representations during training. Pooling thus enables GNNs to scale to large and complex real-world graphs. However, the choice of pooling method strongly influences downstream-applications and task-performance. In fact, the question of which graph properties to preserve during pooling [44, 32, 34], just as the question on the nature and quality of graph datasets in graph learning [13], remains an ongoing debate. It is thus crucial to design expressive, efficient, and interpretable pooling layers that are capable of reliably encoding task-relevant information while reducing the size of input graphs. Most graph pooling literature takes a node-centric view [32]. However, this focus on node-centric rather than edge-centric pooling often leads to the loss of important structural information. Common pooling methods either drop nodes or optimise for a node clustering while treating graph connectivities as a secondary objective. As visualised in Figure 1 and further explored in our work, this frequently leads to counter-intuitive pooling decisions that fail to retain key geometric structures in the graph.
Addressing these concerns, topological and geometric descriptors of graphs are uniquely poised to interoperate structural information into graph pooling. Throughout this work, we treat graphs as metric spaces and assess their geometry via diffusion distances, which naturally work alongside message passing to effectively encode key graph structures. Motivated by ongoing research on novel geometric invariants, we find that generalised measures of size and diversity are especially promising candidates for guiding graph pooling. In particular, we use the magnitude of a metric space, which measures a graph’s structural diversity, to control the loss of structural information during edge pooling [30]. Our work is motivated by successful applications of magnitude across a range of machine learning tasks, such as the evaluation of diversity [31] for generative modelling, boundary detection for images [2], and the study of generalisation for neural networks [3, 4]. Building up on this research, we are the first to propose the use of magnitude in the context of graph learning. We further advance on existing applications by investigating an alternative and closely-related measure, known as the spread of a metric space, to greatly improve the computational efficiency of our methods.
Our main contributions are as follows:
  • We propose MagEdgePool and SpreadEdgePool, novel edge-contraction based pooling layers, which preserve graphs’ structural diversity.
  • We propose the spread of a metric space as a faster and more stable alternative to magnitude and ensure that our algorithms can be computed efficiently.
@@
  • We demonstrate that our pooling methods consistently perform well at graph classification, surpassing alternative pooling layers across a wide range of datasets and pooling ratios.
@@ 2.1 Graph Pooling for Graph Neural Networks
Figure 1: Examples of pooled graphs compared across different pooling layers. Our proposed methods, MagEdgePool and SpreadEdgePool, respect the original graphs’ geometry during pooling. Alternative approaches tend to obscure adjacency relationships to varying extents by creating counter-intuitive edges (Graclus, NMF), disconnecting entire portions of the graphs (TopK, SAGPool), or returning dense representations that do not preserve any geometric structure (DiffPool, MinCut).
@@
As an ongoing field of interest for graph learning, a wide range of pooling methods has been proposed. The most successful of these methods often use a hierarchical approach to sequentially coarsen the graph representation while reducing its size [32]. Research has traditionally focused on node-based pooling via node clustering or node dropping. Amongst node drop pooling methods, Node Decimal Pooling (NDP) [9] is non-trainable, while TopK [21, 11] and SAGPool [27], are trainable approaches. However, all these methods inevitably lose information because they do not select all vertices but remove entire sections of the graph during pooling [32]. Node clustering approaches are similarly either non-trainable, such as Graclus [15] and Non-Negative Matrix Factorization (NMF) [6], or trainable, such as MinCUT [8] and DiffPool [45], which output dense graph representations. In comparison to node drop pooling, clustering-based methods usually require high memory costs [32]. Overall, trainable methods have the potential to better optimise for a specific objective, but might overfit to the task at hand, especially for smaller datasets. By contrast, non-trainable methods can act as a stronger inductive bias on the underlying graph representation, and do not introduce additional trainable parameters or optimisation objectives into the GNN [25]. Edge-based pooling methods have been studied less extensively than node-centric pooling [16, 26] despite their promising potential for encoding connectivities in a more faithful and natural manner. EdgePool [16], the most successful edge-based method, uses iterative edge-contraction based on edge scores, which are learned from features of adjacent nodes. However, EdgePool always pools graphs to half their size, making it less flexible than more adaptive methods. Further, it does not explicitly consider a graph’s topology, beyond posing a constraint on not contracting adjacent edges. Thus, selecting edges based on node features can lead to counter-intuitive decisions during pooling, for example by retaining local structures in strongly-connected communities even at high pooling ratios [16].
Interpretability remains a leading concern [32], and non-trainable approaches are capable of being well-performing and traceable baseline pooling methods [25]. We therefore find that there is strong potential for designing geometry-aware edge pooling operations that make interpretable decisions on which aspects of the graph to retain. Figure 1 further illustrates how many of the aforementioned pooling methods inherently fail to preserve key graph structures even for simple toy examples. Trainable methods in this overview were optimised for a spectral loss following Grattarola et al. [25], but this does not ensure interpretable preservation of graphs’ underlying geometry. Addressing these shortcomings, it is of interest to leverage alternative tools from computational geometry, which allow us to quantify the qualitative difference between graphs. While pooling based on spectral properties has been investigated extensively [9, 25], alternative geometric invariants like curvature or persistent homology have only been explored more recently [44, 20] and have not yet been applied to edge pooling specifically.
@@ 2.2 The Magnitude and Spread of Metric Spaces
We consider an undirected finite graph $G=(X,E)$ with $n$ nodes as a finite metric space consisting of the node set $X$ equipped with a metric $d\colon X\times X\rightarrow\mathbb{R}_{\geq 0}$. Its similarity matrix $\zeta_{X}\subseteq\mathbb{R}^{n\times n}$ is defined by $\zeta_{X}(x,y)=e^{-d(x,y)}$ for $x,y\in X$. This allows us to introduce similarity-dependent notions of the diversity of metric spaces. To this end, we define a weighting on the metric space $(X,d)$, which is a vector $w\in\mathbb{R}^{n}$ such that $\zeta_{X}w=\mathds{1}$, where $\mathds{1}$ is the column vector of ones. Whenever such a weighting exists, the magnitude of the metric space $(X,d)$ is uniquely defined by $\text{Mag}(X)=\sum_{i=1}^{n}w(i)$. This is guaranteed if $\zeta_{X}$ is positive definite, which essentially means that $\zeta_{X}$ is invertible. A finite metric space with positive definite similarity matrix is called positive definite [33]. Metric spaces of negative type are positive definite [30]; this includes $\mathbb{R}^{n}$ equipped with the Euclidean distance [28], effective resistance distances[14], and diffusion distances [12].
Throughout this paper, we will refer to the magnitude of a graph $G$ as the magnitude of its associated metric space $(X,d)$. That is, we define the magnitude of a graph as
[…]
Closely related to magnitude, the spread of a metric space is another measure of ‘size’ introduced by Willerton [41]. Given a metric graph $G=(X,E)$ with the graph metric $d$, its spread is defined by
@@
As diversity measures on graphs, both magnitude and spread summarise the number of distinct sub-communities in a network based on the distance metric and degree of similarity between nodes. This view on structural diversity naturally aligns with our goal of contracting redundant graph structures during pooling. Throughout this work, we investigate to what extent spread is a valid alternative to magnitude in practice. Magnitude is well-studied and has received increasing interest in both theory [30] and application [31]. However, computing it either requires inverting a matrix, solving a system of linear equations [31], or approximating [5], which can be notoriously expensive or unstable to compute. Metric-space spread in comparison can be computed given any distance (obviating the requirement of metric spaces of negative type), making it much more versatile [41]. Moreover, as the sum of reciprocal mean similarities, spread can be calculated or approximated [18] much more efficiently than magnitude and does not require inverting a matrix. Although spread has been studied less extensively [41, 18], there are strong reasons to assume that it shares the same advantages as magnitude. In fact, for a positive definite metric space $X$, we have $\mathrm{Sp}(X)\leq\mathrm{Mag}(X)$ [41, Theorem 2]. Moreover, magnitude and spread coincide for finite homogeneous metric spaces [41, Theorem 3], such as the ring graph in Figure 1. In practice, as we further explore in Section D.1, the magnitude and spread of graphs from real-world datasets, such as NCI1, exhibit nigh-perfect correlation when computed based on diffusion distances, underlining the strong connection between the two quantities.
@@ 2.3 Diffusion Geometry on Graphs
In this work, we use diffusion distances to compute magnitude and spread on graphs. This is motivated by their desirable theoretical properties and the capability of diffusion to aid and act along message passing in GNNs [23]. Diffusion operators are closely associated to random walks and are very efficient at identifying important structures in complex geometries while preserving local and non-linear structures [23]. The key idea is that the eigenvectors of the Markov matrices can be thought of as coordinates for the underlying graph structure [12]. This provides a vector space representation of the graph that can assess the dissimilarity between nodes.
We now briefly detail the type of diffusion distance used throughout this paper [12]. Consider a graph $G=(X,E)$ and its adjacency matrix $A$. The latter could be an affinity matrix derived from a kernel. Let $D$ be the diagonal degree matrix whose diagonal entries $D_{ii}=\sum_{j=0}^{n}A_{ij}$ equal the degree of each vertex. The symmetrically normalised graph Laplacian is defined as $\hat{L}=D^{-\frac{1}{2}}(D-A)D^{-\frac{1}{2}}$. This can be regarded as the simplest example of a Markov transition matrix and represents the probability of moving from one vertex to another. Since $\hat{L}$ is symmetric and positive definite it has positive eigenvalues $1=\lambda_{0}>\lambda_{1}>\lambda_{2}>\cdots\lambda_{N-1}\geq 0$ with eigenvectors $\{\psi_{l}\}_{l}$. This provides a natural embedding of the graph $G$ in Euclidean space given by:
@@
The diffusion distance is defined by :
@@
Proof. See Section B.1.
As a consequence of Theorem 1, the magnitude of any metric graph equipped with this diffusion distance is well defined. Leveraging diffusion distances for our methods has further benefits. The normalised graph Laplacian, which works with the relative connectivity between nodes, is robust to varying node degrees and graph sizes. This ensures that diffusion distances are on comparable scales across graphs, which enables us to compute and compare magnitude and spread directly.
@@ 3.1 Magnitude-Guided Graph Pooling
At the heart of our approach, we use magnitude or spread to monitor and control structural changes in the graph during edge contraction pooling. Edge contraction is chosen as a pooling operation because it respects graph connectivity, while outputting a sparsely-connected graph representation. Conceptually, edge pooling thus aligns well with the goal of making minimal changes to the graph and keeping its diversity and geometry as unchanged as possible. Intuitively, pooled graphs with comparable magnitude will be similar in terms of effective size. That is, the effective number of distinct communities in the graphs are deemed similar based on their diffusion distance i.e. the information-flow between vertices.
Figure 2: Illustrating our proposed pooling method, MagEdgePool, on a graph from the ENZYME dataset across varying pooling ratios. Each edge is coloured by its magnitude difference, which measures the impact its contraction would have on the graph’s structural diversity. Edges with low magnitude differences are most redundant for the graph’s geometry and are collapsed first.
More formally, let $G=(V,E)$ be a graph and denote by $G/{e}$ the graph resulting from the contraction of the edge $e\in E$ in $G$. We choose a pooling ratio $r\in(0,1]$ and aim to reduce the graph to the corresponding number of nodes i.e. $k=\lfloor r\cdot|V|\rceil$. Initially, we set the pooled graph $G^{\prime}=(V^{\prime},E^{\prime}):=G$. To assess which edges to contract first, we determine their importance for the graph’s global geometry by computing a selection score for each edge defined as
@@
That is, we calculate the difference in magnitude between the original graph and the graph for which the edge has been collapsed. In this manner, we score an edge’s relevance for the graph structure by the impact its collapse would have on the graph’s magnitude. At each iteration $i$, we then select the edge with the lowest magnitude difference
@@
and assign $G^{\prime}=G^{\prime}/e_{i}$ where $E_{c}\subseteq E^{\prime}$ is the set of all edges that are not adjacent to an already contracted edge. The edge to contract $e_{i}$ is chosen randomly amongst the possible choices whenever there is more than one valid option. If all edges that meet this requirement have been collapsed, but the pooling ratio is not reached, we re-compute the edge scores on the new graph and repeat the same procedure. We stop if the pooling ratio is reached, i.e. $|V^{\prime}|=\lfloor r\cdot|V|\rceil$, or if there are no edges left in the reduced graph, i.e. $E^{\prime}=\emptyset$. This approach allows us to flexibly reduce graphs to any desired size.
Edge pooling then gives a hard assignment of nodes, where each vertex is assigned to a single super-node in the output graph based on the neighbours it has been merged with. To compress the node features, we average the features of any node that contributed to a pooled super-node. This ensures that information on all nodes’ features is preserved during pooling. Restricting the number of times a vertex can be merged is enforced to prevent our methods from collapsing entire portions of a graph early in the pooling process. This enables a more uniform pooling across the graph, which aids feature preservation and expressivity.
Our methods assume that the most redundant edges will be those whose removal would change the diversity of the graph the least. Magnitude informs us about the global importance of an edge. Intuitively, that means we want to start with merging edges from well connected communities whenever their contraction does not notably change the graph’s effective size. As illustrated in Figure 2, this ensures that globally influential edges will be collapsed late in the pooling progress. For example, the edge which bridges the two parts of this enzyme graph, is scored more highly and thus merged later than well-connected cliques. Redundant local structures are collapsed first, and the overall diffusion geometry is respected. For the graph in Figure 2 this ensures that key topological characteristics, such as the cycle, are retained across the pooling process. Our algorithm is explained in more detail in Section C.4. Whenever we use magnitude to compute the edge scores, we denote our algorithm by MagEdgePool. Similarly, the use of spread is denoted by SpreadEdgePool.

### v2

@@ Abstract
Graph Neural Networks (GNNs) have shown significant success for graph-based tasks. Motivated by the prevalence of large datasets in real-world applications, pooling layers are crucial components of GNNs. By reducing the size of input graphs, pooling enables faster training and potentially better generalisation. However, existing pooling operations often optimise for the learning task at the expense of discarding fundamental graph structures, thus reducing interpretability. This leads to unreliable performance across dataset types, downstream tasks and pooling ratios. Addressing these concerns, we propose novel graph pooling layers for structure-aware pooling via edge collapses. Our methods leverage diffusion geometry and iteratively reduce a graph’s size while preserving both its metric structure and its structural diversity. We guide pooling using magnitude, an isometry-invariant diversity measure, which permits us to control the fidelity of the pooling process. Further, we use the spread of a metric space as a faster and more stable alternative ensuring computational efficiency. Empirical results demonstrate that our methods (i) achieve top performance compared to alternative pooling layers across a range of diverse graph classification tasks, (ii) preserve key spectral properties of the input graphs, and (iii) retain high accuracy across varying pooling ratios.
@@ 1 Introduction
Graph pooling layers are important components of GNN architectures. They are implemented alongside convolutional layers to reduce the size of graph representations during training. Pooling thus enables GNNs to scale to large and complex real-world graphs while regularising the resulting representation. However, the choice of pooling method strongly influences downstream-applications and task-performance. In fact, the question of which graph properties to preserve during pooling, just as the question on the nature and quality of graph datasets in graph learning [15], remains an ongoing debate [55, 41, 44]. It is thus crucial to design expressive, efficient, and interpretable pooling layers that are capable of reliably encoding task-relevant information while reducing the size of input graphs. Most graph pooling literature takes a node-centric view [41]. However, this focus on node-centric rather than edge-centric pooling often leads to the loss of important structural information. Common pooling methods either drop nodes or optimise for a node clustering while treating graph connectivities as a secondary objective. As visualised in Figure˜1 and further explored in our work, this frequently leads to counter-intuitive pooling decisions that fail to retain key geometric structures in a graph. Addressing these concerns, topological and geometric descriptors of graphs are uniquely poised to interoperate structural information into graph pooling.
Throughout this work, we treat graphs as metric spaces and assess their geometry via diffusion distances, which naturally work alongside message passing to effectively encode key graph structures. Motivated by ongoing research on novel geometric invariants, we find that generalised measures of size and diversity are especially promising candidates for guiding graph pooling. In particular, we use the magnitude of a metric space [39], which measures a graph’s structural diversity, to control the loss of structural information during edge pooling. Our work is motivated by successful applications of magnitude across a range of machine learning tasks, such as the evaluation of diversity [40] for latent spaces, boundary detection for images [2], and the study of the generalisation behaviour of neural networks [3, 4]. Building up on this research, we are the first to propose the use of magnitude in the context of graph learning. We further advance on existing applications by investigating an alternative and closely-related measure, known as the spread of a metric space, to substantially improve the computational efficiency of our methods. Our main contributions are as follows:
  • We propose MagEdgePool and SpreadEdgePool, two novel edge-contraction based pooling layers, which preserve graphs’ structural diversity.
  • We investigate the spread of a metric space as a faster and more stable alternative to magnitude and ensure that our algorithms can be computed efficiently.
@@
  • We demonstrate that our pooling methods consistently perform well in graph classification tasks, achieving top accuracies among other pooling layers across a wide range of experimental setups.
@@ 2 Background
We briefly provide background information on graph pooling, the magnitude of metric spaces, and diffusion geometry, taking care to point out related work and how it differs from ours.
@@ 2.1 Graph Pooling for Graph Neural Networks
Figure 1: Examples of graphs pooled to approximately half their original size compared across pooling layers. Our proposed methods, MagEdgePool and SpreadEdgePool, respect the original graphs’ geometry during pooling. Alternative approaches tend to obscure adjacency relationships to varying extents by creating counter-intuitive edges (Graclus, NMF), disconnecting entire portions of the graphs (TopK, SAGPool), or returning dense representations that do not preserve any geometric structure (DiffPool, MinCut).
As an ongoing field of interest for graph learning, a wide range of pooling methods has been proposed, which can be divided into global and hierarchical approaches. Global and hierarchical pooling methods fulfil fundamentally different but complementary roles as distinct components of GNN architectures [32]. Global or flat pooling methods generate graph-level representations by reducing each graph to a single node and are commonly used as readout operations. Examples include mean-pool or sum-pool, which average or sum all node features, SOPool [51], which uses second-order feature information, or DKEPool [13], which learns node distributions. However, global pooling methods discard all topological information, which can reduce the expressivity of these models and decrease their performance [34, 8, 32]. By contrast, hierarchical pooling methods sequentially coarsen graph representations while reducing their size, and are frequently used in alternation with intermediate convolutional or message passing layers [41]. By reducing both features and adjacencies, hierarchical methods have the ability to preserve graphs’ inherent geometry allowing GNNs to learn across multiple coarsening levels.
@@
Research has traditionally focused on node-based hierarchical pooling via node clustering or node dropping. Amongst node drop pooling methods, Node Decimal Pooling (NDP) [10] is non-trainable, while TopK [28, 12] and SAGPool [36], are trainable approaches. However, all these methods inevitably lose information because they do not select all vertices but remove entire sections of the graph during pooling [41]. Node clustering approaches are similarly either non-trainable, such as Graclus [19] and Non-Negative Matrix Factorization (NMF) [6], or trainable, such as MinCUT [9] and DiffPool [56], which output dense graph representations. In comparison to node drop pooling, clustering-based methods usually require high memory costs [41]. Overall, trainable methods have the potential to better optimise for a specific objective, but might overfit to the task at hand, especially for smaller datasets. By contrast, non-trainable methods can act as a stronger inductive bias on the underlying graph representation, and do not introduce additional trainable parameters or optimisation objectives into the GNN [32]. Edge-based pooling methods have been studied less extensively than node-centric pooling [20, 35] despite their promising potential for encoding connectivities in a more faithful and natural manner. EdgePool [20], the most successful edge-based method, uses iterative edge-contraction based on edge scores, which are learned from features of adjacent nodes. However, EdgePool always pools graphs to half their size, making it less flexible than more adaptive methods. Moreover, learning the edge scores during training is computationally expensive [8] necessitating the development of faster and more efficient edge pooling approaches. Further, EdgePool does not explicitly consider a graph’s topology, beyond posing a constraint on not contracting adjacent edges. Thus, selecting edges based on node features can lead to counter-intuitive decisions during pooling, for example by retaining local structures in strongly-connected communities even at high pooling ratios [20]. Interpretability remains a leading concern [41], and non-trainable approaches are capable of being well-performing and traceable baseline pooling methods [32]. We therefore find that there is strong potential for designing geometry-aware edge pooling operations that make interpretable decisions on which aspects of the graph to retain. Figure˜1 further illustrates how many of the aforementioned pooling methods inherently fail to preserve key graph structures even for simple toy examples. Trainable methods in this overview were optimised for a spectral loss following Grattarola et al. [32], but this does not ensure interpretable preservation of graphs’ underlying geometry. Addressing these shortcomings, it is of interest to leverage alternative tools from computational geometry, which allow us to quantify the qualitative difference between graphs. While pooling based on spectral properties has been investigated extensively [10, 32], alternative geometric invariants like curvature or persistent homology have only been explored more recently [55, 25], and have not yet been applied to edge pooling specifically.
@@ 2.2 The Magnitude and Spread of Metric Spaces
Throughout this paper, we consider an undirected finite graph $G=(X,E)$ with $n$ nodes as a finite metric space, consisting of the node set $X$ equipped with a metric $d\colon X\times X\rightarrow\mathbb{R}_{\geq 0}$. Its similarity matrix $\zeta_{X}\subseteq\mathbb{R}^{n\times n}$ is defined by $\zeta_{X}(x,y)=e^{-d(x,y)}$ for $x,y\in X$. This allows us to introduce similarity-dependent notions of the diversity of metric spaces. To this end, we define a weighting on the metric space $(X,d)$, which is a vector $w\in\mathbb{R}^{n}$ such that $\zeta_{X}w=\mathds{1}$, where $\mathds{1}$ is the column vector of ones. Whenever such a weighting exists, the magnitude of the metric space $(X,d)$ is uniquely defined by $\text{Mag}(X)=\sum_{i=1}^{n}w(i)$. This is guaranteed if $\zeta_{X}$ is positive definite, which essentially means that $\zeta_{X}$ is invertible. A finite metric space with positive definite similarity matrix is called positive definite [43]. Metric spaces of negative type are positive definite [39]; this includes $\mathbb{R}^{n}$ equipped with the Euclidean distance [37], effective resistance distances[18], and diffusion distances [14]. Subsequently, we will refer to the magnitude of a graph $G$ as the magnitude of its associated metric space $(X,d)$. That is, we define the magnitude of a graph as
$\text{Mag(G)}=\sum_{x,y\in X}\zeta_{X}^{-1}(x,y).$ (1)
Closely related to magnitude, the spread of a metric space is another measure of ‘size’ introduced by Willerton [52]. Given a metric graph $G$ with the graph metric $d$, its spread is defined by
@@
As diversity measures on graphs, both magnitude and spread summarise the number of distinct sub-communities in a network based on the distance metric and degree of similarity between nodes. This view on structural diversity naturally aligns with our goal of contracting redundant graph structures during pooling. Throughout this work, we investigate to what extent spread is a valid alternative to magnitude. This is motivated by the fact that computing magnitude in practice either requires inverting a matrix, solving a system of linear equations [40], or resort to approximations [5], which can be computationally expensive and numerically unstable. Metric-space spread in comparison can be computed given any distance (obviating the requirement of metric spaces of negative type), making it much more versatile [52]. Moreover, as the sum of reciprocal mean similarities, spread can be calculated or approximated [23] much more efficiently than magnitude and does not require inverting a matrix. Although spread has been studied less extensively, there are strong reasons to assume that it shares the same advantages as magnitude. In fact, for a positive definite metric space $X$, we have $\mathrm{Sp}(X)\leq\mathrm{Mag}(X)$ [52, Theorem 2]. Moreover, magnitude and spread coincide for finite homogeneous metric spaces [52, Theorem 3], such as the ring graph in Figure˜1. In practice, as we further explore in Section˜D.1, the magnitude and spread of graphs from real-world datasets, such as NCI1, exhibit nigh-perfect correlation when computed based on diffusion distances, underlining the strong connection between the two quantities.
@@ 2.3 Diffusion Geometry on Graphs
In this work, we use diffusion distances to compute magnitude and spread on graphs. This is motivated by their desirable theoretical properties and the capability of diffusion to aid and act along message passing in GNNs: Diffusion operators are closely associated to random walks and are efficient at identifying important structures in complex geometries while preserving local and non-linear structures [30]. The key idea is that the eigenvectors of the Markov matrices can be thought of as coordinates for the underlying graph structure [14]. This provides a vector-space representation of the graph that can be used to assess the dissimilarity between nodes.
We now briefly detail the type of diffusion distance used throughout this paper. Consider a graph $G$ and its adjacency matrix $A$. Let $D$ be the diagonal degree matrix whose diagonal entries $D_{ii}=\sum_{j=1}^{n}A_{ij}$ equal the degree of each vertex. The symmetrically normalised adjacency matrix $\hat{A}:=D^{-\frac{1}{2}}AD^{\frac{1}{2}}$ is a Markov transition matrix and represents the probability of moving from one vertex to another. The normalised graph Laplacian is defined as $\hat{L}=I-\hat{A}=D^{-\frac{1}{2}}(D-A)D^{-\frac{1}{2}}$. Since $\hat{L}$ is symmetric and positive definite, it has positive eigenvalues $2\geq\lambda_{0}>\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n-1}\geq 0$ with eigenvectors $\{\psi_{l}\}_{l}$. This provides a natural embedding of the graph $G$ in Euclidean space given by:
@@
The diffusion distance is then defined by the $l^{2}$-norm, i.e.,
@@
As a consequence of ˜1, the magnitude of any metric graph equipped with this diffusion distance is well defined. Leveraging diffusion distances for our methods has further benefits. The normalised graph Laplacian, which works with the relative connectivity between nodes, is robust to varying node degrees and graph sizes. This ensures that diffusion distances are on comparable scales across graphs, which enables us to compute and compare magnitude and spread directly. We note, however, that our pooling methods are flexible and can be applied to a wider range of alternative distances or similarities between nodes, which can be tailored to an application domain.
@@ 3 Methods
We first describe our magnitude-guided graph pooling methods in Section˜3.1 while providing an in-depth theoretical analysis in Section˜3.2.
@@ 3.1 Magnitude-Guided Graph Pooling
At the heart of our approach, we use magnitude or spread to monitor and control structural changes in the graph during edge contraction pooling. Edge contraction is chosen as a pooling operation because it respects graph connectivity, while outputting a sparsely-connected graph representation. Conceptually, edge pooling thus aligns well with the goal of making minimal changes to the graph and keeping its diversity and geometry as unchanged as possible. Intuitively, pooled graphs with comparable magnitude will be similar in terms of effective size. That is, the effective number of distinct communities in the graphs are deemed similar based on their diffusion distance, i.e. based on the information flow between vertices.
Figure 2: Illustrating our proposed pooling method, MagEdgePool, on a graph from the ENZYMES dataset across varying pooling ratios. Each edge is coloured by its magnitude difference, which measures the impact its contraction would have on the graph’s structural diversity. Edges with low magnitude differences are most redundant for the graph’s geometry and are collapsed first.
Formally, let $G=(X,E)$ be a graph and denote by G/eG/{e} the graph resulting from the contraction of the edge $e\in E$ in $G$. We choose a pooling ratio $r\in(0,1]$ and aim to reduce the graph to the corresponding number of nodes i.e. $k=\lfloor r\cdot|X|\rceil$ where $\lfloor{\,}\rceil$ denotes rounding to the nearest integer. Initially, we set the pooled graph $G^{\prime}=(X^{\prime},E^{\prime}):=G$. To assess which edges to contract first, we determine their importance for the graph’s global geometry by computing a selection score for each edge, which we define as
@@
That is, we calculate the difference in magnitude between the original graph and the graph for which the edge has been collapsed. In this manner, we score an edge’s relevance for the graph structure by the impact its collapse would have on the graph’s magnitude. In each iteration $i$, we then select the edge with the lowest magnitude difference
@@
and assign $G^{\prime}=G^{\prime}/e_{i}$ where $E_{c}\subseteq E^{\prime}$ is the set of all edges that are adjacent to an already contracted edge. The edge $e_{i}$ that we contract is selected at random whenever there is more than one valid option. If all edges that meet this requirement have been collapsed, but the pooling ratio is not reached, we re-compute the edge scores on the new graph and repeat this procedure. We stop if the pooling ratio is reached, i.e. $|X^{\prime}|=\lfloor r\cdot|X|\rceil$, or if there are no edges left in the reduced graph, i.e. $E^{\prime}=\emptyset$. This approach allows us to flexibly reduce graphs to any desired size.
Edge pooling then gives a hard assignment of nodes, where each vertex is assigned to a single super-node in the output graph based on the neighbours it has been merged with. To compress the node features, we average the features of any node that contributed to a pooled super-node. This ensures that information on all nodes’ features is preserved during pooling. Note that the feature aggregation function could easily be modified to use sum pooling instead of mean pooling, for example. Restricting the number of times a vertex can be merged is enforced to prevent our methods from collapsing entire portions of a graph early in the pooling process. This enables a more uniform pooling across the graph, which aids feature preservation and expressivity.
In a nutshell: Our methods assume that the most redundant edges will be those whose removal would change the diversity of the graph the least. Magnitude informs us about the global importance of an edge. This implies that we want to start by merging edges from well-connected communities whenever their contraction does not notably change the graph’s effective size.
As illustrated in Figure˜2, our pooling method ensures that globally influential edges will be collapsed late in the pooling progress. For example, the edge which bridges the two parts of this enzyme graph is scored more highly and thus merged later than well-connected cliques. Redundant local structures are collapsed first, and the overall (diffusion) geometry is respected. For the graph in Figure˜2, this ensures that characteristic features of the enzyme, such as the cycle, are preserved across the pooling process. We provide a pseudocode implementation of our algorithm in Section˜C.4. Notice that we may use, mutatis mutandis, spread in lieu of magnitude. Whenever we use magnitude to compute the edge scores, we denote our algorithm by MagEdgePool, else we use the moniker SpreadEdgePool.